OpenStudy (anonymous):

what is the sum of 5/16 inch and 1/4 inch?

jhonyy9 (jhonyy9):

9/16 inch

OpenStudy (anonymous):

r u sure?

OpenStudy (anonymous):

9/16

jhonyy9 (jhonyy9):

yes

OpenStudy (anonymous):

ok thank you

jhonyy9 (jhonyy9):

yw

OpenStudy (anonymous):

convert both fractions to 16th's 1/4 = 4/16 so 5/16 + 4/16 = 9/16